derive an expression for escape and orbital velocity
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
mv^2/r = GMm/r^2Centripetal force = gravitational forceOrbital velocity, for a small mass m orbiting a large mass M:
v = sqrt (GM /r)

Escape velocity: Potential energy relative to infinity = -GMm/r. Escape velocity is the velocity such that the initial KE is just enough to get you to infinity due to conservation of energy So
(1/2)mv^2 = GMm/r.
 
v = sqrt (2GM/r)
 
 
Regards
Arun (askIITians forum expert)